Karlottalotta:

The goal is to increase the power. Therefore, it is necessary to (2 points) decrease the work being done and decrease the time in which the work was completed decrease the work being done and increase the time in which the work is completed increase the work being done or decrease the time in which the work is completed increase the work being done and increase the time in which the work was completed

AZ:

power = work/time if you want to increase the value of power, there are two ways you can manipulate the fraction to get that you either increase the numerator, or you decrease the denominator (or you do both at the same time) To understand what I mean Let's say you had a fraction 1/100 That's equivalent to 0.01 If we increased the numerator (that's the number on top which is 1) 2/100 = 0.02 20/100 = 0.2 200/100 = 2 Do you see how if we increase the numerator, we are increasing the value of the fraction Now if we decrease denominator when we have 1/100 1/50 = 0.02 1/10 = 0.1 1/1 = 1 1/0.1 = 10 Do you see how when we decrease the denominator, we're increasing the values? We went from 0.1 and as we keep decreasing the denominator, we get a bigger number.

AZ:

So for your formula Power = work/time so what do you want to do to work and what do you want to do to the time to increase the power?